CHAPTER 6

QUESTION 1:

Periodic Inventory Using FIFO, LIFO, and Weighted Average Cost Method The units of an item available for sale during the year were as follows Jan. 1 Aug. 13 Purchase 260 units at $342 Nov. 30 Purchase40 units at $357 Inventory 20 units at $360 $7,200 88,920 14,280 $110,400 Available for sale 320 units There are 57 units of the item in the physical inventory at December 31. The periodic inventory system is used. Determine the inventory cost using the (a) the first-in, first-out (FIFO) method; (b) the last-in, first-out (LIFO) method; and (c) the weighted average cost method a. First-in, first-out (FIFO) method b. Last-in, first-out (LIFO) method c. Weighted average cost method

QUESTION 2:

QUESTION 3:

QUESTION 4:

QUESTION 5:
Answer #1

Q1) Calculate ending inventory cost :

Under FIFO = (40*357+17*342) = 20094

Under LIFO = (20*360+37*342) = 19854

Under weighted average method = (110400/320*57) = 19665

Note : Please post each question individually for answer these all question are independent quesstion